OpenStudy (anonymous):

ΔABC is a right triangle where m∠B = 90°. The coordinates of A and B are (5, 0) and (2, 5), respectively. If the x-coordinate of point C is 7, what is the y-coordinate of C ? A. 5 B. 7 C. 8 D. 10

ganeshie8 (ganeshie8):

you're given : A = (5, 0) B = (2, 5)

ganeshie8 (ganeshie8):

say C = (7, y)

OpenStudy (anonymous):

but how do you find the y coordinate?

ganeshie8 (ganeshie8):

use below : product of slopes of perpendicular lines = -1

ganeshie8 (ganeshie8):

find the slope of AB find the slope of BC

ganeshie8 (ganeshie8):

set the product equal to -1 and solve y

ganeshie8 (ganeshie8):

can you find the slopes ?

OpenStudy (anonymous):

i knew how to do it before but i haven't done it in so long so i forgot how loll

ganeshie8 (ganeshie8):

easy, use the slope formula http://www.algebra-class.com/image-files/slope-formula-2.gif

ganeshie8 (ganeshie8):

A = (5, 0) B = (2, 5) slope = ?

OpenStudy (anonymous):

i got 5/-3 after i subtracted them. I think i did it wrong

ganeshie8 (ganeshie8):

Correct !

ganeshie8 (ganeshie8):

next, find the slope of BC

ganeshie8 (ganeshie8):

B = (2, 5) C = (7, y) slope = ?

OpenStudy (anonymous):

im assuming you set the product of y=-1. So using that in the formula I got -6/5 for the slope

ganeshie8 (ganeshie8):

that comes later, first we need to find the slope in terms of y : B = (2, 5) C = (7, y) slope = (y-5)/(7-2) = (y-5)/5

ganeshie8 (ganeshie8):

now set the product of both slopes equal to -1 : (5/-3)((y-5)/5) = -1 solve y

OpenStudy (anonymous):

im thinking the slope is 8. Idk if i solved for y correctly or not

ganeshie8 (ganeshie8):

5 cancels out and you will get : (y-5)/-3 = -1 y - 5 = 3 y= 8

ganeshie8 (ganeshie8):

good job !!!

OpenStudy (anonymous):

lml i got it right ! :D

OpenStudy (anonymous):

thank you ganeshie ur the besttt

ganeshie8 (ganeshie8):

no, you're the best !! yw :)